# There Was Never a Bottleneck in Concept Bottleneck Models

### v1

@@ 5.1 Do CBMs and MCBMs Leak Information?
CelebA CUB
CBM $\bm{92.2\pm 0.7}$ $96.5\pm 0.1$
CEM $92.1\pm 0.8$ $96.5\pm 0.1$
MCBM (low $\gamma$) $92.0\pm 0.8$ $\bm{96.7\pm 0.1}$
MCBM (medium $\gamma$) $92.1\pm 1.0$ $96.4\pm 0.1$
MCBM (high $\gamma$) $91.8\pm 1.1$ $95.8\pm 0.4$
One might suspect that removing nuisance information from the representation could inadvertently eliminate information about the concepts as well. To investigate this, we report concept prediction accuracy for CelebA and CUB in Table 4 (similar trends were observed on the remaining datasets). We observe no significant overall differences across models: while MCBMs show slightly lower concept accuracy on CelebA, they outperform the baselines on CUB for lower $\gamma$. A regular trend, however, is that increasing the hyperparameter $\gamma$ in MCBMs leads to a gradual decrease in concept accuracy, which is consistent with our expectations—stronger regularization may inadvertently suppress features that correlate with the concepts. These results suggest that MCBMs are effective at removing nuisance information while largely preserving the concept-relevant content in the representation.
@@ 5.2 Do MCBMs Yield More Interpretable Representations?
One of the key properties often attributed to CBMs is that their internal representations align with human-interpretable concepts. While this generally holds, we show that MCBMs yield representations that are even more interpretable. To support this claim, we employ two metrics: (i) Centered Kernel Alignment(CKA) [50, 51, 52] , computed between the learned representations and the concept labels (encoded as one-hot vectors), which quantifies the similarity between representation spaces using kernel methods; and (ii) Disentanglement[36], which assesses whether each dimension of the representation $z_{j}$ captures at most one concept $c_{j}$. We argue that these metrics offer meaningful evidence of interpretability because (i) alignment between the representations and the concepts indicates whether the information is organized in a concept-aware fashion; and (ii) disentanglement reflects the extent to which individual concepts are independently encoded in the learned representations. In Figure 5, we observe that: (i) although the representations learned by CBMs and CEMs exhibit substantially higher alignment with the concepts than VMs, their disentanglement levels are not reliably superior; and (ii) MCBMs consistently achieve improvements in both alignment and disentanglement. Thus, we conclude that removing nuisances from the representations leads to more interpretable representations.
5.3 What Are the Pitfalls of Binarizing Multiclass Concepts?
(a) One-vs-Rest
(b) Multiclass
Figure 6: Class Boundaries
(c) One-vs-Rest
(d) Multiclass
Figure 7: Highest likelihood
(a) One-vs-Rest
(b) Multiclass
Figure 8: Second highest likelihood
As discussed in Section 4.1, CBMs require binarizing multiclass concepts to support interventions—i.e., they adopt a One-vs-Rest classification scheme. In contrast, as discussed in Section 4.4, MCBMs provide a natural and principled framework for modeling and intervening on multiclass concepts directly. Binarizing multiclass variables can introduce several issues, particularly in the presence of class imbalance, which is common in real-world applications. To illustrate this limitation, we design an experiment using a four-class spiral dataset with imbalanced class distributions. We train two CBMs for comparison: (i) one that binarizes concepts using a One-vs-Rest approach, and (ii) one that models concepts directly as multiclass variables.
First, as shown in Figure 8, the One-vs-Rest model tends to favor the most frequent class more often than the Multiclass model, particularly when the true class has low prevalence and is similar (i.e., spatially close in the spiral setting) to the dominant class. Second, Figure 8 illustrates that, in the Multiclass model, the predicted class likelihood increases gradually from regions far from the decision boundaries toward the boundaries, resulting in a well-calibrated probability distribution. Such patterns are absent in the One-vs-Rest model. Finally, in Figure 8, we observe that the second-highest predicted likelihood in the One-vs-Rest model is close to one in certain regions—indicating that multiple classes are simultaneously assigned high confidence. In contrast, the Multiclass model assigns low second-highest likelihoods in regions far from the boundaries, with values increasing only near the decision boundaries.
@@ 6 Conclusions
In this paper, we argue that—contrary to common belief—Concept Bottleneck Models (CBMs) do not enforce a true bottleneck: while representations are encouraged to retain concept-related information, they are not constrained to discard nuisance information. This limitation results in a lack of theoretical guarantees for intervention procedures and reduced interpretability in the representation space. To address this, we propose Minimal Concept Bottleneck Models (MCBMs), which introduce an Information Bottleneck in the representation space via an additional loss term derived through variational approximations. Beyond establishing a proper bottleneck, our formulation ensures that interventions do not violate Bayesian principles and introduces greater design flexibility compared to standard CBMs. Through empirical analysis, we show that CBMs—and their existing variants—fail to effectively remove nuisance information from the representation, even when such information is irrelevant to the target task. In contrast, MCBMs successfully eliminate nuisances while preserving concept-relevant information, resulting in more interpretable representations and principled intervention mechanisms. Finally, we highlight a common limitation of CBMs: the need to binarize multiclass concepts due to architectural constraints. MCBMs natively support multiclass concepts, thereby overcoming the drawbacks introduced by this binarization process
@@ Appendix A Complete Graphical Models
Figure 9: Graphical models of the different systems described. We consider $m$ concepts and $m$-dimensional representations. Inputs $\bm{x}$ are defined by some concepts $\{c_{j}\}_{j=1}^{m}$ and nuisances $\bm{n}$; and targets $\bm{y}$ are defined by $\bm{x}$ (gray arrows). Vanilla models obtain the representations $\{z_{j}\}_{j=1}^{m}$ from $\bm{x}$ through the encoder $p_{\theta}(z|x)$ (black arrows) and solve the task $\hat{y}$ sequentially through the task head $q_{\phi}(y|z)$ (green arrows). Concept Bottleneck Models try to make a prediction $\hat{c}_{j}$ of each concept $c_{j}$ from one representation $z_{j}$ through the concept head $q(\hat{c}_{j}|z_{j})$ (blue arrows). Minimal CBMs (ours) try to make a prediction $\hat{z}_{j}$ of each representation $z_{j}$ from one concept $c_{j}$ through the representation head $q(\hat{z}_{j}|c_{j})$ (red arrows).
@@ Appendix C Training Algorithm of MCBMs
17:         Sample noise $\epsilon\sim\mathcal{N}(0,I)$triangleright Reparameterization trick with only one sample
@@
19:         Task prediction: $\hat{\bm{y}}^{(k)}\leftarrow g^{y}_{\phi}\left(\bm{z}^{(k)}\right)$triangleright Similar to VMs
@@
22:              Concept $j$ prediction: $\hat{c}^{(k)}_{j}\leftarrow g^{c}_{\phi,j}\left(z^{(k)}_{j}\right)$triangleright Similar to CBMs
@@
24:              Representation $j$ prediction: $\hat{z}^{(k)}_{j}\leftarrow g^{z}_{j}\left(c^{(k)}_{j}\right)$triangleright Novelty in MCBMs
@@
Appendix D Experiments Details
D.1 Hyperparameters for Section 5
@@
CIFAR-10 MPI3D Shapes3D CelebA CUB
$f_{\theta}$ architecture ResNet20 ResNet20 ResNet20 ResNet18 InceptionV3
$f_{\theta}$ pretraining None None None ImageNet ImageNet
$g^{y}_{\phi}$ hidden layers 64 64 64 64 256
low $\gamma$ 1 1 1 0.05 0.1
medium $\gamma$ 5 3 3 0.1 0.3
high $\gamma$ 10 5 5 0.3 0.5
number of epochs 50 50 50 120 250
batch size 128 128 128 128 128
optimizer SGD SGD SGD SGD SGD
learning rate $6\times 10^{-3}$ $6\times 10^{-3}$ $6\times 10^{-3}$ $2\times 10^{-2}$ $2\times 10^{-2}$
momentum 0.9 0.9 0.9 0.9 0.9
weight decay $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$
scheduler Step Step Step Step Step
step size (epochs) 20 20 20 50 100
scheduler $\gamma$ 0.1 0.1 0.1 0.1 0.1
D.2 Datasets
CIFAR-10
The CIFAR-10 dataset consists of 60,000 color images of size 32×32, evenly distributed across 10 object classes, with 50,000 training and 10,000 test samples. Among these, 6 classes correspond to animals and 4 to vehicles. We define the target task $\bm{y}$ as a binary classification between animals and vehicles. The subset $\bm{n}_{y}$ includes one vehicle and one animal class—specifically, airplane and bird—while $\bm{c}$ comprises the remaining 8 object classes.

### v2

@@ 4.1 Do CBMs and MCBMs Leak Information?
CIFAR-10 CUB
CBM $84.8\pm 0.2$ $96.3\pm 0.1$
CEM $84.8\pm 0.2$ $96.3\pm 0.1$
ARCBM $84.3\pm 0.2$ $96.2\pm 0.1$
SCBM $84.3\pm 0.2$ $\bm{96.5\pm 0.1}$
MCBM (l $\gamma$) $\bm{84.9\pm 0.1}$ $96.3\pm 0.2$
MCBM (m $\gamma$) $\bm{84.9\pm 0.2}$ $96.1\pm 0.1$
MCBM (h $\gamma$) $84.8\pm 0.1$ $95.8\pm 0.3$
One might worry that removing nuisance information could also inadvertently eliminate information about the concepts. To test this, we report concept prediction accuracy for CIFAR-10 and CUB in Table 5 (note that all models reach 100% accuracy on MPI3D and Shapes3D). We can observe that (i) no model consistently outperforms the others—some preserve more concept information in CIFAR-10, while others perform slightly better in CUB; and (ii) increasing $\gamma$ in MCBMs gradually reduces concept accuracy, as stronger regularization may suppress features correlated with the concepts. These results indicate that MCBMs effectively remove nuisance information while largely preserving concept-relevant content.
How does this affect task performance?
Table 6: Task accuracy
MPI3D Shapes3D CIFAR-10 CUB
Vanilla $\pm 0.0 100.0±0.010$$\pm 0.0$ $2.1\pm 0.6 77$$7.4\pm 0.3$
CBM $99.9\pm 0.0$ $100.0\pm 0.0$ $3.8\pm 0.1 77$$7.6\pm 0.2$
CEM $\pm 0.0 100.0±0.010$$\pm 0.0$ $3.1\pm 0.3 77$$7.5\pm 0.3$
ARCBM $24.2\pm 0.6$ $29.7\pm 0.4$ $68.9\pm 0.3$ $75.5\pm 0.8$
SCBM $24.2\pm 0.5$ $29.7\pm 0.3$ $62.3\pm 0.0$ $73.5\pm 0.4$
MCBM (l $\gamma$) $92.7\pm 1.1$ $100.0\pm 0.0$ $72.4\pm 0.5$ $78.3\pm 0.6$
MCBM (m $\gamma$) $46.0\pm 0.6$ $32.2\pm 1.5$ $0.8\pm 0.2 77$$7.4\pm 0.4$
MCBM (h $\gamma$) $24.9\pm 0.3$ $30.1\pm 0.2$ $70.5\pm 0.6$ $73.5\pm 1.5$
As previously discussed, when $\bm{n}_{y}$ is non-empty, restricting $\bm{z}$ to contain only information about $\bm{c}$ should reduce performance on $\bm{y}$: if $\bm{c}$ alone is insufficient to solve $\bm{y}$, then $\bm{z}$ cannot achieve perfect accuracy. In Table 6, we observe the following: (i) CBMs and CEMs reach task accuracy comparable to (or even higher than) VMs, indicating that they also rely on $\bm{n}_{y}$ to predict $\bm{y}$; (ii) ARCBMs and SCBMs achieve lower task accuracy, as they impose a bottleneck after the representations (see Appendix D); and (iii) MCBMs show decreasing task accuracy as $\gamma$ increases, reflecting the stricter bottleneck applied to the representations. Importantly, this reduction in task accuracy should not be viewed negatively: since all models achieve similar concept accuracy (see Table 5), it indicates that predictions rely less on nuisance information.
@@ 4.2 Do MCBMs Yield More Interpretable Representations?
One of the key properties often attributed to CBMs is that their internal representations align with human-interpretable concepts. While this generally holds, we show that MCBMs yield representations that are even more interpretable. To support this claim, we employ two metrics: (i) Centered Kernel Alignment(CKA) [38, 39, 40], which measures the similarity between learned representations and concept labels (encoded as one-hot vectors); and (ii) Disentanglement[41], which assesses whether each dimension of the representation $z_{j}$ captures at most one concept $c_{j}$. These metrics capture interpretability from complementary perspectives: (i) alignment indicates whether the information is organized in a concept-aware fashion; and (ii) disentanglement reflects the extent to which individual concepts are independently encoded in the representations. As shown in Figure 4: (i) CBMs, CEMs, ARCBMs, and HCBMs achieve stronger alignment with concepts than Vanilla Models, yet they do not consistently improve disentanglement; and (ii) MCBMs consistently improve both alignment and disentanglement, particularly as $\gamma$ increases, reflecting the removal of additional nuisances. These results suggest that explicitly eliminating nuisances leads to more interpretable representations.
4.3 Do MCBMs Enable More Reliable Interventions?
(a) CIFAR-10
(b) CUB
Figure 5: Error (y-axis) versus percentage of concepts intervened (x-axis) across different models.
Another key property often attributed to CBMs is their capacity to support interventions. Although standard CBM intervention methods suffer from theoretical limitations (see Section 5), they can still be applied in practice. Accordingly, for CBMs we follow the procedure in Section 2.3, for ARCBMs and SCBMs we use the procedure described in Appendix D, and for MCBMs we apply the method in Section 2.4. To evaluate intervention reliability across models, we adopt the standardized protocol of Koh et al. [9], which tracks how prediction error changes as the number of intervened concepts increases. Following Shin et al. [42], we intervene on the concepts predicted with the lowest confidence. Results for CIFAR-10 and CUB are shown in Figure 5 (note that interventions have minimal effect on MPI3D and Shapes3D, since—as shown in Table 6—the concepts are not strong predictors of the task). From these results, four main observations emerge: (i) CBMs may even increase error when multiple concepts are intervened—an effect of nuisance information leaking into the representation due to the absence of a proper bottleneck; (ii) ARCBMs and SCBMs mitigate this by applying an Information Bottleneck after the representations, which improves intervention effectiveness while leaving interpretability unchanged (Figure 4); (iii) in MCBMs, intervention gains remain largely invariant to $\gamma$, as indicated by the nearly parallel trends across its different values; and (iv) at low and medium $\gamma$, MCBMs deliver the strongest intervention performance, especially when fewer than 100% of the concepts are intervened, clearly outperforming ARCBMs and SCBMs.
5 Other Fundamental Theoretical Flaws of CBMs
As briefly discussed in Section 2.3, CBMs—unlike MCBMs—do not provide a principled mechanism to estimate $p(z_{j}\mid c_{j})$. To enable interventions, two assumptions are typically introduced: (i) multiclass concepts are handled using a One-vs-Rest scheme, and (ii) interventions are implemented via the sigmoid inverse function, i.e., $p(z_{j}\mid c_{j})\approx\delta\left(z_{j}-\sigma^{-1}(c_{j})\right)$. These assumptions are not only ad hoc but also theoretically incorrect, as we explain and illustrate with toy experiments below.
One-vs-Rest Limitations
Figure 6: CBM (top) versus MCBM (bottom): class boundaries (left), highest predicted likelihood (middle), and second-highest predicted likelihood (right).
One-vs-Rest strategies exhibit several limitations: (i) individual binary classifiers tend to be biased toward the negative class, and (ii) the predicted probabilities across classifiers are typically uncalibrated [43]. To illustrate these issues, we design an experiment where the concepts are defined based on a four-class spiral dataset with imbalanced class distributions. We train: (i) a CBM with One-vs-Rest binarization, and (ii) an MCBM modeling concepts directly as multiclass variables.
The results reveal three major shortcomings of CBMs trained with the One-vs-Rest strategy. First, Figure 6 (left) shows that they overpredict the most frequent class, especially when the true class is rare and spatially close to the dominant. Second, Figure 6 (middle) shows that CBMs produce poorly calibrated predictions, lacking the smooth likelihood transitions observed in MCBMs. Finally, Figure 6 (right) shows that CBMs often assign near-one likelihoods to multiple classes simultaneously, whereas MCBMs confine high secondary likelihoods to regions of overlap, producing more reliable uncertainty estimates. Importantly, these effects may not be captured by standard metrics like accuracy, yet they represent fundamental flaws from a probabilistic perspective.
Sigmoid Inverse Function
Figure 7: Density (y-axis) vs. $z$ (x-axis)
[…]
@@ 6 Conclusions
In this paper, we argue that—contrary to common belief—Concept Bottleneck Models (CBMs) do not enforce a true bottleneck: although representations are encouraged to retain concept-related information, they are not constrained to discard nuisance information. This limitation undermines interpretability and provides no theoretical guarantees for intervention procedures. To address this, we propose Minimal Concept Bottleneck Models (MCBMs), which introduce an Information Bottleneck in the representation space via an additional loss term derived through variational approximations. Beyond enforcing a proper bottleneck, our formulation ensures that interventions remain consistent with Bayesian principles. Empirically, we show that CBMs and their variants fail to remove non-concept information from the representation, even when irrelevant to the target task. In contrast, MCBMs effectively eliminate such information while preserving concept-relevant content, yielding more interpretable representations and principled interventions. Finally, we highlight fundamental flaws in the intervention process of CBMs, which MCBMs overcome due to their principled design.
@@ Appendix A Complete Graphical Models
Figure 8: Graphical models of the different systems described. We consider $m$ concepts and $m$-dimensional representations. Inputs $\bm{x}$ are defined by some concepts $\{c_{j}\}_{j=1}^{m}$ and nuisances $\bm{n}$; and targets $\bm{y}$ are defined by $\bm{x}$ (gray arrows). Vanilla models obtain the representations $\{z_{j}\}_{j=1}^{m}$ from $\bm{x}$ through the encoder $\color[rgb]{0,1,1}p_{\theta}(\bm{z}|\bm{x})$ (cyan arrows) and solve the task $\hat{\bm{y}}$ sequentially through the task head $\color[rgb]{0,1,0}q_{\phi}(\hat{\bm{y}}|\bm{z})$ (green arrows). Concept Bottleneck Models make a prediction $\hat{c}_{j}$ of each concept $c_{j}$ from one representation $z_{j}$ through the concept head $\color[rgb]{0,0.49609375,1}q(\hat{c}_{j}|z_{j})$ (blue arrows). Minimal CBMs make a prediction $z_{j}$ of each representation $z_{j}$ from one concept $c_{j}$ through the representation head $\color[rgb]{1,0,0}q(z_{j}|c_{j})$ (red arrows).
@@ Appendix C Training Algorithm of MCBMs
17:   Sample noise $\epsilon\sim\mathcal{N}(0,I)$ $\triangleright$ Reparameterization trick with only one sample
@@
19:   Task prediction: $\hat{\bm{y}}^{(k)}\leftarrow g^{y}_{\phi}\left(\bm{z}^{(k)}\right)$ $\triangleright$ Similar to VMs
@@
22:     Concept $j$ prediction: $\hat{c}^{(k)}_{j}\leftarrow g^{c}_{\phi,j}\left(z^{(k)}_{j}\right) \triangleright$ Similar to CBMs
@@
24:     Representation $j$ prediction: $\hat{z}^{(k)}_{j}\leftarrow g^{z}_{j}\left(c^{(k)}_{j}\right) \triangleright$ Novelty in MCBMs
@@
Appendix D Hard Concept Bottleneck Models
As discussed in Section 3, one of the most successful approaches to mitigating information leakage is the family of Hard Concept Bottleneck Models (HCBMs) [18]. Unlike MCBMs, which enforce an Information Bottleneck (IB) directly at the representation level $\bm{z}$, HCBMs operate by pruning the information contained in $\bm{z}$ before producing the task prediction $\hat{\bm{y}}$. In doing so, they encourage $\hat{\bm{y}}$ to depend solely on $\bm{c}$, thereby yielding interventions that are more reliable in practice. This approach differs fundamentally from MCBMs: whereas Equation 4 in MCBMs minimizes over $Z_{j}$, thereby removing nuisances directly from the representation $z_{j}$, HCBMs optimize at the prediction level, focusing on $\hat{Y}$. Formally, they are trained to minimize:
$\vskip 2.0pt\min_{\hat{Y}}I(\hat{Y};X|C)=\min_{\theta,\phi}\mathbb{E}_{p(x,c_{j})}\left[D_{KL}\left(p_{\theta,\phi}(\hat{\bm{y}}|x)||p(\hat{\bm{y}}|c)\right)\right]$ (24)
That is, the optimization is performed directly over the predictions $\hat{\bm{y}}$. To implement this, it defines:
$p_{\theta,\phi}(\hat{\bm{y}}|x)=\iiint q_{\phi}\left(\hat{\bm{y}}|\hat{\bm{c}}^{b}\right){\color[rgb]{1,0,1}p\left(\hat{\bm{c}}^{b}|\hat{\bm{c}}\right)}q\left(\hat{\bm{c}}|\bm{z}\right)p_{\theta}\left(\bm{z}|\bm{x}\right)\,d\hat{\bm{c}}^{b}\,d\hat{\bm{c}}\,d\bm{z}$ (25)
where $q_{\phi}\left(\hat{\bm{y}}\mid\hat{\bm{c}}^{b}\right)$ denotes the new task head, $q\left(\hat{\bm{c}}\mid\bm{z}\right)$ the concept head, and $p_{\theta}\left(\bm{z}\mid\bm{x}\right)$ the encoder, as defined in Section 2. The distribution
$p\left(\hat{c}_{j}^{b}|\hat{c}_{j}\right)=\delta\left(\hat{c}_{j}^{b}-\Theta\left(\hat{c}_{j}-0.5\right)\right)$ (26)
referred to as the binarizing head, applies the Heaviside step function $\Theta$ to produce a binary version $\hat{c}^{b}_{j}\in\{0,1\}$ of $\hat{c}_{j}$. In this way, HCBMs enforce an ad-hoc Information Bottleneck by predicting $\hat{\bm{y}}$ from binarized concept representations. This process is schematized in Figure 9.
How are Interventions Performed in HCBMs?
Because HCBMs introduce a bottleneck immediately before predicting $\hat{\bm{y}}$, the intervention process is more straightforward than in standard CBMs. More specifically, interventions are carried out according to:
$p(\hat{\bm{y}}|c_{j}=\alpha,\bm{x})=\iint q_{\phi}(\hat{\bm{y}}|\hat{c}^{b}_{j},\bm{\hat{c}}^{b}_{\backslash j})p(\hat{c}^{b}_{j}|c_{j}=\alpha)p_{\theta}(\bm{\hat{c}}^{b}_{\backslash j}|\bm{x})\,d\hat{c}^{b}_{j}\,d\bm{\hat{c}}^{b}_{\backslash j}$ (27)
Here, $p(\hat{c}^{b}_{j}|c_{j}=\alpha)$ is approximated as $\delta(\hat{c}^{b}_{j}-\alpha)$, while the other distributions are available in closed form. Although this procedure provides stronger guarantees than the intervention mechanisms in standard CBMs, two main issues remain:
  1. (i)
    The optimization is performed over the predictions $\hat{\bm{y}}$ instead of the representation $\bm{z}$. As a result, the representations themselves are not necessarily interpretable, as evidenced in Figure 4, which limits one of the core motivations for adopting CBMs in the first place.
  2. (ii)
    HCBMs still require binarizing multiclass concepts, which introduces theoretical limitations and practical drawbacks, as further discussed in Section 5.
Figure 9: Graphical models of HCBMs with two concepts and two-dimensional representations. Hard Concept Bottleneck Models obtain a binarized version $\hat{c}^{b}_{j}$ of each predicted concept $\hat{c}_{j}$ through the binarizing head $\color[rgb]{1,0,1}p(\hat{c}^{b}_{j}\mid\hat{c}j)$ (fuchsia arrows). Unlike the models in Figure 2, HCBMs predict the task output $\hat{\bm{y}}$ from the binarized concepts $\hat{c}^{b}_{j}$ using the new task head $\color[rgb]{0,1,0}q_{\phi}(\hat{\bm{y}}\mid\bm{z})$ (green arrows).
Appendix E Experiments Details
E.1 Hyperparameters for Section 4
@@
MPI3D Shapes3D CIFAR-10 CUB
$f_{\theta}$ architecture ResNet20 ResNet20 2 conv. layers InceptionV3
$f_{\theta}$ pretraining None None None ImageNet
$g^{y}_{\phi}$ hidden layers 64 64 64 256
low $\gamma$ 1 1 0.1 0.05
medium $\gamma$ 3 3 0.3 0.1
high $\gamma$ 5 5 0.5 0.3
number of epochs 50 50 200 250
batch size 128 128 128 128
optimizer SGD SGD Adam SGD
learning rate $6\times 10^{-3}$ $6\times 10^{-3}$ $1\times 10^{-4}$ $2\times 10^{-2}$
momentum 0.9 0.9 0. 0.9
weight decay $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$
scheduler Step Step Step Step
step size (epochs) 20 20 80 100
scheduler $\gamma$ 0.1 0.1 0.1 0.1
E.2 Datasets